# Polarons and Exciton-Polarons in Two-Dimensional Polar Materials

### v1

@@ Abstract
We propose a macroscopic theory of optical phonons, Fröhlich polarons, and exciton-polarons in two-dimensional (2D) polar crystalline monolayers. Our theory extends the classical macroscopic formulation of the electron-phonon problem in three-dimensional (3D) polar crystals to the new generation of 2D materials. Similarly to the 3D case, in our approach, the effective electron-phonon Hamiltonian is parametrized solely in terms of macroscopic experimentally accessible quantities – 2D polarizabilities of the monolayer at low and high frequencies. We derive the dispersion of long wave length longitudinal optical (LO) phonons, which can be viewed as a 2D form of the Lyddane-Sachs-Teller relation, and study the formation of 2D Fröhlich polarons by adopting the intermediate coupling approximation. Finally, we apply this approach to excitons in polar 2D crystals and derive an effective potential of the electron-hole interaction dressed by LO phonons. Due to the specific dispersion of LO phonons, the polarons and exciton-polarons in 2D materials exhibit unique features not found in their 3D counterparts. As an illustration, the polaron and exciton-polaron binding energies are computed for a representative set of 2D polar crystals, demonstrating the interplay between dimensionality, polarizability of materials, and electron-phonon coupling.
Polarons, quasiparticles arising from the interaction of charge carriers with lattice vibrations, have played a pivotal role in understanding charge transport and optical properties in polar materials [1]. The concept of polaron was initially formulated as an electron interacting with the induced displacement field in dielectric media, the so-called Landau-Pekar model [2, 3, 4] in which polaronic effects result in autolocalization of the electron in a selfconsistent potential well. Alternatively, a weak and intermediate coupling approach implies a field theoretical picture of the electron dressed by a cloud of longitudinal optical (LO) phonons, known as the Fröhlich polaron [5, 6, 7]. Polaronic effects in bulk polar crystals such as alkali halides and oxides are well studied both theoretically and experimentally [8, 9], providing important benchmarks for strong-coupling electron-phonon physics. Recent advances in this direction include accurate diagrammatic Monte-Carlo results [10, 11], and first-principle calculations [12, 13, 14].
@@
Despite rapid progress in the fabrication and study of two-dimensional (2D) materials [35], there are a handful of observations of polaronic effects in 2D. This includes hexagonal boron nitride (hBN) [36], MoS2 monolayer [37], SrTiO3 surface [38, 39]. On the theory side, most studies of 2D polarons in different physical situations adopt an ad hoc replacement of bare 3D to 2D Coulomb potential in the electron-phonon Hamiltonian [40, 41, 11, 42, 43]. Recently, a first principle extension of Landau-Pekar polaron theory, which accounts for the proper 2D screening, has been developed [44]. Very recently, the exciton-polaron problem for a single layer GaN was addressed within many-body theory [45, 46] by perturbatively correcting the kernel in the Bethe-Salpeter equation [47].
@@
Figure 1: (a) Schematic of a 2D polar crystal layer hosting an electron-hole pair dressed by LO phonons, leading to the formation of exciton-polaron. (b) Dispersion of LO and TO phonon modes for a hBN monolayer. Here the dielectric constant of surrounding media is set $\varepsilon=1$. The remaining parameters are presented in Table 4.
Phonon dispersion. We consider 2D optical phonon modes in a polar monolayer located at the plane $z=0$. Let us define the mass-weighted displacement field $\bm{\xi}=\sqrt{nM}\mathbf{S}$, where $\mathbf{S}=\mathbf{U}_{+}-\mathbf{U}_{-}$ is the relative displacement of positively and negatively charged ions, $n=N/\Sigma$ is the sheet density of unit cells, and $M=m_{+}m_{-}/(m_{+}+m_{-})$ is the reduced ionic mass. The Lagrangian density of the system reads
@@
where the coefficients are uniquely determined from the high and low frequency asymptotic values of the macroscopic 2D polarization $\mathbf{P}=\partial W^{\rm 2D}/\partial\mathbf{E}$. Specifically, here $\omega_{\rm t}$ is the transverse optical (TO) phonon frequency, $\alpha_{\infty}$ is the high-frequency 2D polarizability, and $\gamma=\omega_{\rm t}\sqrt{\alpha_{0}-\alpha_{\infty}}$ where $\alpha_{0}$ is the static polarizability, see Supplemental Material (SM) for details. Apparently, the difference $\alpha_{0}-\alpha_{\infty}$ characterizes the ionic contribution to the polarizability of the monolayer.
The Euler-Lagrange equations, generated by Eq. 1, consist of the equation of motion (EOM) for the 2D displacement $\bm{\xi}(\mathbf{r})$ and the 3D Poisson equation for the electric potential $\varphi(\mathbf{r},z)$. By considering plane-wave solutions, $\bm{\xi}(\mathbf{r})\sim e^{i\mathbf{k}\cdot\mathbf{r}}$, $\varphi(\mathbf{r},z)\sim e^{i(\mathbf{k}\cdot\mathbf{r}+k_{z}z)}$, and defining the in-plane electric potential in the $\mathbf{k}$-space as $\varphi_{\mathbf{k}}^{\rm 2D}=\int\varphi_{\mathbf{k},k_{z}}{\rm d}k_{z}/(2\pi)$, we find
@@
where we introduced the 2D screening lengths as $r_{\infty}=\alpha_{\infty}/(2\varepsilon_{0})$, $r_{0}=\alpha_{0}/(2\varepsilon_{0})$. Equation (4) perfectly agrees the dispersion obtained from heuristic arguments supported by direct first principal calculations in Ref. [52] and recently observed experimentally [56]. Remarkably, in our approach, as expected physically in the long wave length limit, the relation between frequencies of TO an LO phonons involves only experimentally accessible macroscopic polarizabilities. The LO-TO degeneracy at $k=0$ as well as a nontrivial dispersion of LO mode are the manifestations of the dielectric screening in 2D. It is worth noting that in many instances, the quantity $\varepsilon+\alpha k/2\varepsilon_{0}$ enters as an effective 2D dielectric constant, and therefore Eq. (4) can be viewed as a 2D form of the Lyddane-Sachs-Teller relation. We visualize the LO phonon dispersion for hBN monolayer, as depicted in Fig. 1 (b) 222 For the sake of simplicity in all explicit calculations we assume freestanding atomic monolayers, i.e. $\varepsilon=1$. The remaining parameters are summarized in Table 4..
@@
Table 1: Material parameters taken from experimental and first-principal calculations data for a set of 2D polar materials: TO phonon energy $\hbar\omega_{\rm t}$, screening lengths $r_{0}$, $r_{\infty}$444the screening lengths are extracted from Ref. [44] as $r_{0[\infty]}=\epsilon_{0[\infty]}d/2$, where ϵ0[]subscriptitalic-$\epsilon_{0[\infty]}$ is the static [high-frequency] dielectric constant, $d$ is the monolayer thickness., electron and hole effective masses $m_{e}$, $m_{h}$, the bandgap $E_{g}$ (asterisk marks the direct-gap materials). Calculated renormalized masses $m_{e}^{*}=m_{e}(1+\alpha_{e})$, $m_{h}^{*}=m_{h}(1+\alpha_{h})$, polaron binding energies $E^{\rm b}_{e}$, $E^{\rm b}_{h}$, bare exciton binding energy $E_{\rm X}^{0}$, and exciton-polaron binding energy $E_{\rm X}^{\rm b}$. The values of $r_{0}$, $r_{\infty}$, $m_{e}$, $m_{h}$ are taken from Ref. [44].
Material $\hbar\omega_{\rm t}$ (meV) $r_{0}$ (nm) $r_{\infty}$ (nm) $m_{e}/m_{e}^{*}$ $m_{h}/m_{h}^{*}$ $E^{\rm b}_{e}$ (meV) $E^{\rm b}_{h}$ (meV) $E_{\rm X}^{0}/E_{\rm X}^{\rm b}$ (meV) $E_{g}$ (eV)
h-BN $172.3$[52] $1.076$ [52] $0.780$ [52] $0.83$$/1.00$ $0.65$$/0.77$ $129.3$ $116.3$ $1617/1869$ $4.69^{*}$[44]
HfS2 $17.81$[58] $13.98$ $2.974$ $0.24$$/0.54$ $0.44$$/1.07$ $123.2$ $155.1$ $197.8/444.6$ $1.31$ [58]
HfSe2 $11.10$[59] $21.75$ $4.246$ $0.18$$/0.43$ $0.23$$/0.57$ $89.56$ $98.51$ $127.5/314.9$ $0.64$ [59]
GaN $73.30$[60] $1.109$ $0.756$ $0.24$$/0.29$ $1.35$$/2.02$ $55.16$ $132.0$ $1304/1591$ $6.32^{*}$[60]
AlN $74.15$[61] $0.763$ $0.466$ $0.51$$/0.71$ $1.49$$/2.53$ $105.4$ $186.5$ $2064/2891$ $2.81$ [62]
ZrS2 $29.82$[63] $10.62$ $2.826$ $0.31$$/0.58$ $0.26$$/0.48$ $142.9$ $133.7$ $238.5/428.6$ $1.25$ [64]
Figure 2: Screening length ratio dependence of (a) $I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$, and (b) $I_{m}(\sigma_{0},\sigma_{e,{\rm t}})$ characterizing the polaron binding energy and mass renormalization, respectively. The polar effects become more pronounced with the increase of parameter $\sigma_{0}$, resulting in the enhancement of both $I_{\rm b}$, and $I_{m}$. The reduction of $\sigma_{e,{\rm t}}$ corresponds to increase of phonon energy and [or] free electron effective mass, both resulting in the increase of polaron binding and mass renormalization. Asterisks in panel (a) illustrate the representative polar atomic monolayers.
@@
Fröhlich Hamiltonian and polaron problem. The electric potential of Eq. (3), generated by the ionic displacement, acts on the electrons, generating electron-phonon coupling. Formally, the electron-phonon Hamiltonian is derived by inserting the functional $\varphi[\bm{\xi}_{\mathbf{k}}]$ into the Lagrangian (1) and canonically quantizing the displacement field, see SM. We stress that the last term in Eq. (1) is essential to restore the correct Hamiltonian of LO phonons. The final electron-phonon Hamiltonian reads
@@
$\hat{H}_{e-{\rm ph}}=\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}+\sum_{\mathbf{k}}\hbar\omega_{{\rm l},k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\\ +\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}^{\dagger}_{\mathbf{k}}\right),$ (5)
@@
$V_{k}=ie\omega_{\rm t}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k},$ (6)
which depends on the LO phonon dispersion and contains a typical 2D screening factor.
The Hamiltonian of Eq. (5) contains three length scales: $r_{0}$, $r_{\infty}$, and $r_{i,{\rm t}}=\sqrt{\hbar/(2\omega_{\rm t}m_{i})}$, $i=e\,[h]$ for electron [hole]. The two ratios, $\sigma_{0}=r_{0}/r_{\infty}$ and $\sigma_{i,{\rm t}}=\varepsilon r_{i,{\rm t}}/r_{\infty}$, define two independent parameters, characterizing the 2D Fröhlich polaron problem, which should be contrasted to the 3D problem determined by a single dimensionless coupling constant.
We analyze the properties of 2D polaron by adopting the LLP intermediate coupling approach [6], which is expected to be accurate in most realistic situations. Firstly, we note the conservation of total momentum, $[\hat{\mathbf{p}}_{e}+\sum_{\mathbf{k}}\hbar\mathbf{k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}},\hat{H}_{e-{\rm ph}}]=0$, and eliminate the electron coordinate by performing a unitary transformation with the operator $\hat{S}=\exp\left(i\left(\mathbf{p}/h-\sum_{\mathbf{k}}\mathbf{k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\right)\cdot\mathbf{r}_{e}\right)$. Next, the phonon field is shifted with $\hat{U}=\exp\left(\sum_{\mathbf{k}}\left(f(\mathbf{k})\hat{a}^{\dagger}_{\mathbf{k}}-f^{*}(\mathbf{k})\hat{a}_{\mathbf{k}}\right)\right)$, acting as $\hat{U}^{-1}\hat{a}_{\mathbf{k}}\hat{U}=\hat{a}_{\mathbf{k}}+f(\mathbf{k})$, where $f(\mathbf{k})$ is a $C$-number serving as a variational parameter. The variational anzats reads $|\psi\rangle=\hat{U}|\psi_{0}\rangle$, assuming $\hat{a}_{\mathbf{k}}|\psi_{0}\rangle=0$. The variational minimization leads to the following expression for polaron energy (see SM for the details):
$E_{e}^{\rm pol}\approx-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\alpha_{e}}.$ (7)
Here we introduce the notations $E_{0}=e^{2}/(4\pi\varepsilon_{0}r_{\infty})$, $\alpha_{e}=E_{0}I_{m}(\sigma_{0},\sigma_{e,{\rm t}})/(\hbar\omega_{\rm t})$, and dimensionless quantities
@@
$I_{m}$ $=\int\limits_{0}^{\infty}\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)x^{3}}{(1+\sigma_{0}x)^{2}\left(1+\sigma_{e,{\rm t}}^{2}x^{2}\sqrt{\frac{1+x}{1+\sigma_{0}x}}\right)^{3}}\,{\rm d}x.$ (9)
The first term in Eq. (7) corresponds to the polaron binding energy $E^{\rm b}_{e}=E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$, and the second term is the polaron kinetic energy with rescaled effective mass $m_{e}^{*}=m_{e}(1+\alpha_{e})$. Fig. 2 (a), (b) illustrate the factors $I_{\rm b}(\sigma_{0},\sigma_{\rm t})$, $I_{m}(\sigma_{0},\sigma_{\rm t})$ characterizing the polaron binding energy and mass renormalization, respectively. The parameter $\sigma_{0}$ characterizes the strength of polar effects. Therefore, the increase of $\sigma_{0}$ enhances both $I_{\rm b}$ and $I_{m}$. The parameter $\sigma_{e,{\rm t}}\propto 1/\sqrt{\omega_{\rm t}m_{e}}$ is the inverse energy scale of phonon energy and effective mass. Hence, the reduction of $\sigma_{e,{\rm t}}$ enhances both $I_{\rm b}$ and $I_{m}$, in line with the case of bulk polarons, where the polaron binding energy scales as $E_{e}^{\rm pol}\propto\sqrt{\omega_{\rm LO}m_{e}}$ [54]. We recall that the actual values of polaron binding energy and mass renormalization include the material-specific factors $E_{0}$ and $\hbar\omega_{\rm t}$, as it follows from Eq. (7). The respective data for several representative materials is shown in Table 4. The obtained values of polaron binding energy essentially differ from that reported in Ref. [44], where the ab-initio generalization of Landau-Pekar model is developed. We attribute this discrepancy to (i) the strong electron-phonon coupling limit, corresponding to Landau-Pekar model [54, 65], and (ii) the trivial limit of 2D LO phonon dispersion, considered in Refs. [53, 44], which strongly affects the electron–phonon interaction matrix element.
Notably, the resulting expressions for the polaron binding energy and mass renormalization are not of analytical form, unlike the case of bulk polarons [6]. However, substituting $\sqrt{(1+x)/(1+\sigma_{0}x)}\approx 1/\sqrt{\sigma_{0}}$, we get
$I_{\rm b}\approx\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left(\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}\right.\\ \left.+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log(\tilde{\sigma_{e,{\rm t}}})}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log(\sigma_{0})}{\sigma_{0}-1}\right),$ (10)
[…]
@@
Figure 3: Effective potential of electron-hole interaction $V_{\rm eff}$ (solid curve) and its approximate analytical expression $V_{\rm eff}^{\rm an}$ (dashed curve) for (a) h-BN, and (b) HfSe2 monolayer. The dotted curves illustrate the KR potentials $V_{\rm KR}(r,r_{\infty})$ and $V_{\rm KR}(r,r_{0})$.
Figure 4: Spatial profile of exciton wave function for (a) h-BN, and (b) HfSe2 monolayer. Here solid orange curves illustrate the exciton-polaron wave function, and black dotted curves correspond to bare exciton state. (c) Map of 2D polar materials illustrating the relation of single polaron and exciton-polaron binding energies.

### v2

@@ Abstract
We propose a macroscopic theory of optical phonons, Fröhlich polarons, and exciton-polarons in two-dimensional (2D) polar crystalline monolayers. Our theory extends the classical macroscopic formulation of the electron-phonon problem in three-dimensional (3D) polar crystals to the new generation of 2D materials. Similarly to the 3D case, in our approach, the effective electron-phonon Hamiltonian is parametrized solely in terms of macroscopic experimentally accessible quantities – 2D polarizabilities of the monolayer at low and high frequencies. We derive the dispersion of long wave length longitudinal optical (LO) phonons, which can be viewed as a 2D form of the Lyddane-Sachs-Teller relation, and study the formation of 2D Fröhlich polarons by adopting the intermediate coupling approximation. Finally, we apply this approach to excitons in polar 2D crystals and derive an effective potential of the electron-hole interaction dressed by LO phonons. Due to a specific dispersion of LO phonons, polarons and exciton-polarons in 2D materials exhibit unique features not found in their 3D counterparts. As an illustration, the polaron and exciton-polaron binding energies are computed for a representative set of 2D polar crystals, demonstrating the interplay between dimensionality, polarizability of materials, and electron-phonon coupling.
Polarons, quasiparticles arising from the interaction of charge carriers with lattice vibrations, have played a pivotal role in understanding transport and optical properties of polar materials [1]. The original concept of Landau-Pekar polaron [2, 3, 4] corresponds to an electron that locally polarizes the ionic crystals, which reduces its energy and may lead to autolocalizaton in the adiabatic self-consistent potential well. An alternative and complementary field theoretical picture of the electron dressed by a cloud of longitudinal optical (LO) phonons is known as the Fröhlich polaron [5, 6, 7]. Polaronic effects in bulk polar crystals such as alkali halides and oxides are well studied both theoretically and experimentally [8, 9], providing important benchmarks for strong-coupling electron-phonon physics. Here recent advances include accurate diagrammatic Monte-Carlo (MC) results [10, 11], and first-principle calculations [12, 13, 14].
@@
Despite rapid progress in the fabrication and study of two-dimensional (2D) materials [35], there are a handful of observations of polaronic effects in 2D. This includes hexagonal boron nitride (hBN) [36], MoS2 monolayer [37], SrTiO3 surface [38, 39]. Theoretically, polarons in 2D have been mostly studied in the context of surface polarons that appear due to interaction with phonons on surfaces of 3D polar crystals [40, 41, 11, 42]. Only recently, the problem of intrinsically 2D polarons has been addressed within a first principle extension of Landau-Pekar polaron theory [43], and perturbatively by computing polaron correction to Landau levels in the presence of magnetic field [44]. The exciton-polaron problem for a single layer GaN was addressed within ab initio many-body theory [45, 46] by perturbatively correcting the kernel in the Bethe-Salpeter equation. Very recently, quasi 2D exciton-polarons were studied with path integral MC [47], adopting a model electron-phonon coupling of Refs. [48, 49].
@@
Figure 1: (a) Schematic picture of a 2D polar crystal layer hosting an electron-hole pair dressed by LO phonons, leading to the formation of exciton-polaron. (b) Dispersion of LO and TO phonon modes for a hBN monolayer.
Phonon dispersion. We consider 2D optical phonon modes in a polar monolayer located at the plane $z=0$. The Lagrangian density of the system in terms of the mass-weighted displacement field $\bm{\xi}$ reads
@@
where the coefficients are uniquely determined from the high and low frequency asymptotic values of the macroscopic 2D polarization $\mathbf{P}=-\partial W^{\rm 2D}/\partial\mathbf{E}$. Specifically, here $\omega_{\rm t}$ is the transverse optical (TO) phonon frequency, $\alpha_{\infty}$ is the high-frequency 2D polarizability, and $\gamma=\omega_{\rm t}\sqrt{\alpha_{0}-\alpha_{\infty}}$ where $\alpha_{0}$ is the static polarizability, see Supplemental Material (SM) Sec. I for details. Apparently, the difference $\alpha_{0}-\alpha_{\infty}$ characterizes the ionic contribution to the polarizability of the monolayer.
The Euler-Lagrange equations, generated by Eq. (1), consist of the equation of motion (EOM) for the 2D displacement $\bm{\xi}(\mathbf{r})$ and the 3D Poisson equation for the electric potential $\varphi(\mathbf{r},z)$. By considering plane-wave solutions, $\bm{\xi}(\mathbf{r})\sim e^{i\mathbf{k}\cdot\mathbf{r}}$, $\varphi(\mathbf{r},z)\sim e^{i(\mathbf{k}\cdot\mathbf{r}+k_{z}z)}$, and defining the in-plane electric potential in the $\mathbf{k}$-space as $\varphi_{\mathbf{k}}^{\rm 2D}=\int\varphi_{\mathbf{k},k_{z}}{\rm d}k_{z}/(2\pi)$, we find
@@
where screening lengths are defined as $r_{\infty}=\alpha_{\infty}/(2\varepsilon_{0})$, $r_{0}=\alpha_{0}/(2\varepsilon_{0})$. Equation (4) perfectly agrees the dispersion obtained from heuristic arguments supported by direct first principal calculations in Ref. [53] and recently observed experimentally [56]. Remarkably, in our approach, as expected physically in the long wave length limit, the relation between frequencies of TO an LO phonons involves only experimentally accessible macroscopic polarizabilities. The LO-TO degeneracy at $k=0$ as well as a nontrivial dispersion of LO mode are the manifestations of the dielectric screening in 2D. It is worth noting that in many instances, the quantity $\varepsilon+\alpha k/2\varepsilon_{0}$ enters as an effective 2D dielectric constant, and therefore Eq. (4) can be viewed as a 2D form of the Lyddane-Sachs-Teller relation. In Fig. 1(b) we visualize the LO phonon dispersion for hBN monolayer 222 For the sake of simplicity in all explicit calculations we assume freestanding atomic monolayers, i.e. $\varepsilon=1$. The remaining parameters are summarized in Table 4..
@@
Table 1: Calculated polaron binding energies $E^{\rm b}_{e}$, $E^{\rm b}_{h}$, statically screened ($E_{\rm X}^{0}$), exciton-polaron ($E_{\rm X}^{\rm b}$), and bare ($E_{\rm X}^{\infty}$) exciton binding energies, renormalized effective masses $m_{e}^{*}=m_{e}(1+\chi_{e})$, $m_{h}^{*}=m_{h}(1+\chi_{h})$. Material parameters taken from experimental and first-principal calculations data for a set of 2D polar materials: TO phonon energy $\hbar\omega_{\rm t}$, screening lengths $r_{0}$, $r_{\infty}$444the screening lengths are extracted from Ref. [43] as $r_{0[\infty]}=\epsilon_{0[\infty]}d/2$, where ϵ0[]subscriptitalic-$\epsilon_{0[\infty]}$ is the static [high-frequency] dielectric constant, $d$ is the monolayer thickness., electron and hole effective masses $m_{e}$, $m_{h}$ (taken from Ref. [43]).
Material $E^{\rm b}_{e}$ (meV) $E^{\rm b}_{h}$ (meV) $E_{\rm X}^{0}/E_{\rm X}^{\rm b}/E_{\rm X}^{\infty}$ (meV) $m_{e}/m_{e}^{*}$ $m_{h}/m_{h}^{*}$ $\hbar\omega_{\rm t}$ (meV) $r_{0}$ (nm) $r_{\infty}$ (nm)
h-BN $129.3$ $116.3$ $1617/1865/2007$ $0.83$$/1.00$ $0.65$$/0.77$ $172.3$[53] $1.076$ [53] $0.780$ [53]
GaN $55.16$ $132.0$ $1304/1586/1660$ $0.24$$/0.29$ $1.35$$/2.02$ $73.30$[58] $1.109$ $0.756$
AlN $105.4$ $186.5$ $2064/2873/2839$ $0.51$$/0.71$ $1.49$$/2.53$ $74.15$[59] $0.763$ $0.466$
HfSe2 $89.56$ $98.51$ $127.5/306.9/421.4$ $0.18$$/0.43$ $0.23$$/0.57$ $11.10$[60] $21.75$ $4.246$
HfS2 $123.2$ $155.1$ $197.8/434.2/615.6$ $0.24$$/0.54$ $0.44$$/1.07$ $17.81$[61] $13.98$ $2.974$
ZrS2 $142.9$ $133.7$ $238.5/420.6/619.2$ $0.31$$/0.58$ $0.26$$/0.48$ $29.82$[62] $10.62$ $2.826$
Figure 2: (a) The dimensionless binding energy, and (b) mass renormalization rate versus the screening length ratios $\sigma_{0}$, $\sigma_{e,{\rm t}}$. (c) The dimensionless binding energy, and (d) mass renormalization rate versus $\sigma_{0}$ for fixed $\sigma_{e,{\rm t}}$, as shown in legend. Asterisks in panels (a), (b) illustrate the representative polar atomic monolayers.
@@
Fröhlich Hamiltonian and polaron problem. The electric potential of Eq. (3), produced by the ionic displacement $\bm{\xi}$, acts on the electrons, generating electron-phonon coupling. Formally, the electron-phonon Hamiltonian is derived by inserting the functional $\varphi[\bm{\xi}_{\mathbf{k}}]$ into the Lagrangian (1) and canonically quantizing the displacement field, see SM, Sec. II, III, IV A. We stress that the last term in Eq. (1) is essential to restore the correct Hamiltonian of LO phonons. The final electron-phonon Hamiltonian reads
@@
$\hat{H}=\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}+\sum_{\mathbf{k}}\left(\hbar\omega_{{\rm l},k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}+\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+{\rm h.c.}\right)\right)$ (5)
@@
$V_{k}=i\omega_{\rm t}\sqrt{\frac{\hbar e^{2}(r_{0}-r_{\infty})}{4A\varepsilon_{0}\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k},$ (6)
which depends on the LO phonon dispersion and contains a typical 2D screening factor. Here $A$ denotes the normalization area.
The Hamiltonian of Eq. (5) contains four length scales: two screening lengths $r_{0}$ and $r_{\infty}$, the oscillator scale $r_{i,{\rm t}}=\sqrt{\frac{\hbar}{2\omega_{\rm t}m_{i}}}$, and the Bohr radius $a_{i,{\rm B}}=\frac{4\pi\varepsilon_{0}\hbar^{2}}{e^{2}m_{i}}$, where $i=e\,[h]$ for electron [hole]. Their independent ratios, e. g., $\sigma_{0}=r_{0}/r_{\infty}$, $\sigma_{i,{\rm t}}=\varepsilon r_{i,{\rm t}}/r_{\infty}$, and $\sigma_{i,{\rm B}}=a_{i,{\rm B}}/r_{\infty}$ define three independent parameters, characterizing the 2D Fröhlich polaron problem, which should be contrasted to the 3D problem determined by a single dimensionless coupling constant. The reason for this complication is a non-locality of 2D screening.
We analyze the properties of 2D polaron by adopting the LLP intermediate coupling approach [6], which is expected to be accurate in most realistic situations. Firstly, we eliminate the electron coordinate by performing a unitary transformation with the operator $\hat{S}=\exp\left\{i\left(\mathbf{p}/h-\sum_{\mathbf{k}}\mathbf{k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\right)\cdot\mathbf{r}_{e}\right\}$. Next, the phonon field is shifted with $\hat{U}=\exp\left\{\sum_{\mathbf{k}}\left(f_{\mathbf{k}}\hat{a}^{\dagger}_{\mathbf{k}}-f^{*}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\right)\right\}$, acting as $\hat{U}^{-1}\hat{a}_{\mathbf{k}}\hat{U}=\hat{a}_{\mathbf{k}}+f_{\mathbf{k}}$, where $f_{\mathbf{k}}$ is a $C$-number serving as a variational parameter. The variational anzats reads $|\psi\rangle=\hat{U}|\psi_{0}\rangle$, assuming $\hat{a}_{\mathbf{k}}|\psi_{0}\rangle=0$. The variational minimization leads to the following expression for polaron energy (see SM Sec. IV B for the details):
$E_{e}^{\rm pol}\approx-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\chi_{e}},$ (7)
where $E_{0}=\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}$, $\chi_{e}=E_{0}I_{m}(\sigma_{0},\sigma_{e,{\rm t}})/\hbar\omega_{\rm t}$, and we defined dimensionless scaling factors
@@
$I_{m}$ $=\int\limits_{0}^{\infty}\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)x^{3}}{(1+\sigma_{0}x)^{2}\left[1+\sigma_{e,{\rm t}}^{2}x^{2}\sqrt{\frac{1+x}{1+\sigma_{0}x}}\right]^{3}}\,{\rm d}x.$ (9)
The first term in Eq. (7) defines the polaron binding energy $E^{\rm b}_{e}=E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$, and the second term its effective mass $m_{e}^{*}=m_{e}(1+\chi_{e})$. Within the LLP scheme, the dependence on the Bohr radius $\sigma_{B}$ trivially factors as $E_{0}$ in Eq. (7). A nontrivial dependence on the other two parameters is encoded in the scaling factors $I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$ and $I_{m}(\sigma_{0},\sigma_{e,{\rm t}})$, which, as shown in Figs. 2 (a), (b), grow monotonically with an increase in the parameter $\sigma_{0}-1\sim\alpha_{0}-\alpha_{\infty}$ that characterizes polarizability of the lattice and the coupling strength. However, in contrast to 3D the observed growth is highly nonlinear for $\sigma_{0}-1>1$, as illustrated in Figs. 2 (c), (d) for several fixed values of $\sigma_{e,{\rm t}}$. The strong deviation from linearity for larger $\sigma_{0}$ comes from the denominators in Eqs. (9), and is a peculiarity of 2D materials. This is remarkably different from bulk polarons, where at the LLP level, the binding energy and mass renormalization scale linearly with polarizability of the lattice [6]. The length scale $r_{e,t}\propto 1/\sqrt{\omega_{\rm t}m_{e}}\propto\sigma_{e,{\rm t}}$ roughly determines the spatial extension of the polaron. This explains the reduction of polaronic effects with the increase of $\sigma_{e,{\rm t}}$, see Fig. 2.
The binding energies and effective masses of polarons, computed from Eqs. (7)-(9) for several representative materials, are shown in Table 4. Unlike the case of 3D polarons [6], in general, the final results involve integrals that are not expressible in elementary functions. However, for realistic not too large values of $\sigma_{e,{\rm t}}$, the following substitution $\sqrt{(1+x)/(1+\sigma_{0}x)}\approx 1/\sqrt{\sigma_{0}}$, in the integrand in Eq. (8) can be used to get a simplified form,
$I_{\rm b}\approx\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left[\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}\right.\\ \left.+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log\tilde{\sigma}_{e,{\rm t}}}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log\sigma_{0}}{\sigma_{0}-1}\right],$ (10)
where $\tilde{\sigma}_{e,{\rm t}}=\sigma_{e,{\rm t}}/\sigma_{0}^{1/4}$. This expression represents an excellent estimate for the polaron binding energy, with a few percent relative error in most realistic situations. Similarly, an approximate analytic form of the mass scaling factor $I_{m}$ can be derived; see SM, Sec. IV B.
The polaron energies in Table 4 are notably different from those obtained in Ref. [43] from a generalized Landau-Pekar model. In principle, this is not surprising because the adiabatic Landau-Pekar approximation is known to underestimate the binding energy for weak and moderate coupling strengths [54, 63]. In particular, our result for the binding energy in h-BN is almost an order of magnitude larger than that in Ref. [43]. Admittedly, both LLP and Landau-Pekar approximations are variational and therefore both provide upper bounds to the exact energy. In addition, even within the formal Landau-Pekar approach, our Hamiltonian Eq. (5) produces results different from the model in [43]. The reason is the presence of the full phonon dispersion in Eq. (6), missing in [49, 43], but becoming important for a moderately strong coupling (see SM, Sec. IV C).
@@
Figure 3: Effective potential of electron-hole interaction $V_{\rm eff}$ (solid curve) and its approximate form $V_{\rm eff}^{\rm an}$ (dashed curve) for (a) h-BN, and (b) HfSe2 monolayers. The dotted curves show the KR potentials $V_{\rm KR}(r,r_{\infty})$ and $V_{\rm KR}(r,r_{0})$.
Figure 4: Map of 2D materials illustrating the relation of polaron and exciton-polaron binding energies. Here $\Delta E^{\rm b}_{\rm X}=E^{\rm\infty}_{\rm X}-E^{\rm b}_{\rm X}$. "*" and "+" stand for electron and hole polaron binding energies, respectively.